Eleni owns 4 shares of a business. The current share price is 36 per share. The company announced the terms of issuing a share capital increase with preferences.

The company will issue new shares of € 20 per share with a ratio of 1 new share to 4 old ones. What is the price of the right?
Answer #1

2 3 No of share Price 36 New price 20 No of right shares New theoretical price Price of right 32.8 10 3.2 11

B D No of share 4 Price 36 New price 20 No of right shares New theoretical price =(D4*D5+D7*D6)/5 Price of right =D5-D9 13